# Estimation of Concept Explanations Should be Uncertainty Aware

### 13 Dec 2023 v1

@@ Appendix F Additional experiment details
Scene labels considered in Section 6.
@@
F.1 Addition results for Section 6
We report also the tau (Wikipedia, 2023b) distance from concept explanations computed by Simple as a measure of explanation quality. Kendall Tau is a standard measure for measuring distance between two ranked lists. It does so my computing number of pairs with reversed order between any two lists. Since Simple can only estimate the importance of concepts that are correctly annotated in the dataset, we restrict the comparison to only over concepts that are attributed non-zero importance by Simple.
@@ F.1 Addition results for Section 6
Appendix G Ablation Study
G.1 Uncertainty of Concept Activations
As explained in Section 3.1, we estimate the uncertainty on concept activations using a measure on predictability of the concept as shown in Proposition 1. In this section we evaluate the quality of estimated uncertainty and compare with other (potentially simpler) variants of uncertainty estimation. More crisply, we ask the following question.
@@ G.1 Uncertainty of Concept Activations
We conduct our study using the ResNet-18 model pretrained on Places365 and Pascal dataset that were discussed in Section 6. We use human-provided concept annotations to train per-concept linear classifier on the representation layer. We retained only 215 concepts of the total 720 concepts that have at least two positive examples in the dataset. We then evaluated the per-concept linear classifier on a held-out test set to obtain macro-averaged accuracy. The concepts with poor accuracy are the ones that cannot be classified linearly using the representations. Therefore the error rate per concept is the ground-truth for uncertainty that we wish to quantify.
We may now evaluate the goodness of uncertainty: $\epsilon$ of U-ACE by comparing it with ground-truth (error-rate); observe they are both K-dimensional vectors. We report two measures of similarity in Table 4: (1) Cosine-Similarity (Cos-Sim) between $\epsilon$ and error-rate, (2) Jaccard Similarity (JS) (https://en.wikipedia.org/wiki/Jaccard˙index) between top-k least uncertain concepts identified using error-rate and $\epsilon$. For any two vectors $u,v$, and their top-k sets: $S_{1}(u),S_{2}(v)$, the Cos-Sim and JS are evaluated as follows.
@@
We will now introduce two other variants of estimating $\epsilon$.
[…]
Distribution Fit. Inspired by ProbCBM proposed in Kim et al. (2023a), we estimate $\epsilon$ from the data as a learnable parameter through distribution fitting. We assume normal distrbution of the noise and model the standard deviation as a linear projection of the feature vector. The model is summarized below.
We obtain the observed score of a concept given an example: $\mathbf{x}$ using the multi-modal model: $g,g_{text}$ and the text description of the $k^{th}$ concept $T_{k}$. The concept score is modeled to be distributed by a normal distribution whose mean and standard deviation are linear functions of the feature representation of the model-to-be-explained: $f^{[-1]}(\mathbf{x})$. We optimize the value for $[\vec{p_{1}},\vec{p_{2}},\dots,\vec{p_{K}}],[\vec{q_{1}},\vec{q_{2}},\dots,\vec{q_{K}}]$ through gradient descent on the objective $\mathcal{\beta}=MLL(\mathcal{D},g)+\beta\times\mathbb{E}_{\mathcal{D}}\mathbb{E}_{k}[KL(\mathcal{N}(0,I)\|\mathcal{N}(\mu_{k}(\mathbf{x}),\sigma_{k}(\mathbf{x})))]$ very similar to the proposal of Kim et al. (2023a). We picked the best value of $\beta$ and obtained $\epsilon^{DF}$ by averaging over all the examples: $\epsilon^{DF}=\mathbb{E}_{\mathcal{D}}[\sigma_{k}(\mathbf{x})]$.
@@
Table 4: Evaluation of uncertainties estimated using U-ACE, MC sampling and Distribution Fit (see text for their description). Cos-Sim is the cosine-similarity with ground-truth value of uncertainty. The next three columns show Jaccard similarity between the top-k concepts ranked by ground-truth uncertainty and each of the three methods.
@@
G.2 Bayesian Estimation and Significance of Prior
@@
We replicate the study on Broden dataset of Section 6 of Table 1 with two new baselines. We replace the linear model estimation of U-ACE described in Section 3 with an out-of-the-box Bayesian Regression estimator available from sklearn [^1] https://scikit-learn.org/stable/modules/generated/sklearn.linear˙model.BayesianRidge.html, which we refer as Bayes Regr. Effectively, Bayes Regr. is different from U-ACE only in the prior. We also compare with the estimation of fitting using Bayes Regr. but when the input is perturbed with the noise estimated by U-ACE. We refer to this baseline as Bayes Regr. with MC.
@@
G.3 Effect of Regularization Strength on Y-CBM and O-CBM
@@
Hyperparameter tuning is tricky for concept explanations since they lack a ground-truth or validation set. The reported results for Y-CBM and O-CBM in the main paper used the default value of the regularization strength of the corresponding estimator, which is Lasso [^2] https://scikit-learn.org/stable/modules/generated/sklearn.linear˙model.Lasso.html for Y-CBM and LogisticRegression [^3] https://scikit-learn.org/stable/modules/generated/sklearn.linear˙model.LogisticRegression.html for O-CBM. Both the estimators are part of sklearn. We had to reduce the default regularization strength of Y-CBM to $\alpha=10^{-3}$ so that estimated weights are not all 0. The $\kappa$ of U-ACE is somewhat arbitrarily set to $0.02$ on Broden dataset for sparse explanation with non-zero weight for only 20-30% of the concepts.
@@
Appendix H Extension of Simulation Study
Concept Y-CBM O-CBM U-ACE
red or blue -75.4 -1.8 0.1
blue or red 21.9 -1.9 0
green or blue -1.4 1.6 0
blue or green -23.1 1.6 0
@@ Appendix I Evaluation using CUB dataset
Wah et al. (2011) released a bird dataset called CUB with 11,788 images and 200 bird species. Moreover, each bird image is annotated with one of 312 binary attributes indicating the presence or absence of a bird feature. Koh et al. (2020) popularized an improved version of the dataset that retained only 112 clean attribute annotations. We evaluate using the cleaner dataset released by Koh et al. (2020) owing to their popularity in evaluating CBMs. We train a pretrained ResNet-18 model using the training split of CUB dataset. We then compute the explanation (i.e. saliency of concepts) using the test split. Similar to the evaluation of Section 6, we quantify the quality of explanations using distance from explanations computed using true concept annotations when using Simple.
@@
Table 8: Distance of top-k salient concepts computed using Simple and different estimation methods shown in the first row (lower the better). TCAV does well overall and U-ACE performs the best among methods without access to concept annotations.
@@
Table 9: Average overlap between top-k salient concepts computed using Simple and different estimation methods shown in the first row (higher the better). TCAV does well overall and U-ACE performs the best among methods without access to concept annotations.

### 05 Apr 2024 v2

@@ G.1 Simulated Study
Scene labels considered in Section 6.1.
@@
G.2 Further details for Imagenet experiment of Section 6.2
Table 6 contains the list of labels we considered for evaluation on the Imagenet dataset of Section 6.2.
Table 7 extends with more results the Table 4 of the main content.
Index Name Known spurious features
1 dogsled snow, dog, tree, trees, husky
2 howler monkey trunk, green, branches, branch, vegetation
3 seat belt, seatbelt passenger, window, sunglasses, van
4 ski tree, trees, snow, mountain, person, sunglasses
5 volleyball sand, players, player, setter, scoreboard, net
6 boathouse lake, water, dock, boat, shore
7 bee flower, daisy, petals
8 plate food, table, dining
9 barracouta, snoek person, face, hands, hand, sunglasses, cap
10 llama hay, grass, green, greens
11 rhinoceros beetle hand, head, palm, person, fingers
12 dowitcher water, reflection, lake, shoal, sandbar
13 white wolf, Arctic wolf net, fence, fencing
14 dragonfly blurry, green, flower, plant
15 gorilla, Gorilla gorilla green, tree, grass, trunk
16 shovel snow
17 doormat, welcome mat door
18 ruddy turnstone, Arenaria interpres mud, shore, sand, land, seashore, beach
19 albatross, mollymawk water, sea, ocean
20 sax, saxophone player, players, playing
21 balance beam, beam player, person, sport, arms, legs
22 bathing cap, swimming cap face, chest, person, swimmer, diver, gymnast
23 puck, hockey puck player, bat, ice, arena
24 dining table, board chairs, chair, corner
25 rugby ball ground, green, players, player
26 dock, dockage, docking facility boat, boats, ship, yacht, water, sea, lake
27 padlock chain, chains, door
28 potter’s wheel hands, hand, person, face, head
29 ping-pong ball player, human, hands, arms, arm
30 paddle, boat paddle human, arm, arms, body, lifevest, water
31 unicycle, monocycle road, body, human, arms, face, arm
32 ice lolly, lolly, lollipop, popsicle mouth, eyes, face, human, head, hand, lip, lips
33 beaver water, lake, waterbody
34 mountain tent mountain, lake, water, hill, snow
35 indri, indris, Indri indri, Indri brevicaudatus sky, tree, leaf, leaves, trunk, vegetation, green
36 seashore, coast, seacoast, sea-coast seawater, ocean, water, sea
37 sunglass cheeks, face, head, person, nose
38 bulbul branch, tree, leaves, sky, leaf
39 alp sky, clouds, blue
40 Arabian camel desert, sand, hot, water, ground
Table 6: Classes selected for evaluation in Section 6.2 and known spurious features in the last column. The known spurious features are obtained from Salient-Imagenet
volley ball Y-CBM setter, ball, airship
O-CBM setter, ball, clay
U-ACE sports, setter, ball
Gorilla Y-CBM animals, siamang, mower
O-CBM chimp, black, mastiff
U-ACE siamang, gibbon, macaque
Rugby Y-CBM rugby, knee, missile
   ball O-CBM rugby, soccer, bulldog
U-ACE rugby, sports, player
hockey Y-CBM hockey, player, nail
   puck O-CBM hockey, shaft, heater
U-ACE hockey, sports, gandola
swimming Y-CBM hat, regions, bandaid
   cap O-CBM swimming, bald, meets
U-ACE swimming, sports, snorkel
Table 7: More results extending the Table 4
G.3 Addition results for Section 6.1
We report also the tau [Wikipedia, 2023b] distance from concept explanations computed by Oracle as a measure of explanation quality in Table 8. Kendall Tau is a standard measure for measuring distance between two ranked lists. It does so my computing number of pairs with reversed order between any two lists. Since Oracle can only estimate the importance of concepts that are correctly annotated in the dataset, we restrict the comparison to only over concepts that are attributed non-zero importance by Oracle.
@@ 40 Arabian camel desert, sand, hot, water, ground
Appendix H Ablation Study
H.1 Uncertainty of Concept Activations
As explained in Section 3.1, we estimate the uncertainty on concept activations using a measure on predictability of the concept as shown in Proposition 1. In this section we evaluate the quality of estimated uncertainty and compare with other (simpler) variants of uncertainty estimation. More crisply, we ask the following question.
@@ H.1 Uncertainty of Concept Activations
We will now introduce two new variants of estimating uncertainty.
Alongside the measure of uncertainty of U-ACE discussed in Section 3.1, which we denote with $\epsilon$, we will present two alternate ways of measuring uncertainty below.
MC Sampling. We may simply repeat the estimation procedure several times (denote by S) with different seed and data split to sample multiple concept activation vectors: $\{a_{k}^{(1)},a_{k}^{(2)},\dots,a_{k}^{(S)}\}\quad k\in[1,K]$. We empirically estimate per-concept uncertainty by averaging over examples: $\epsilon^{MC}=\mathbb{E}_{\mathbf{x}\in\mathcal{D}}[V([\mathbf{x}^{T}a_{k}^{(1)},\mathbf{x}^{T}a_{k}^{(2)},\dots,\mathbf{x}^{T}a_{k}^{(S)}])]$. Where $V(\bullet)$ is the sample variance: $V(b_{1},b_{2},\dots,b_{S})=\frac{\sum_{s}(b_{s}-\frac{\sum_{s}b_{s}}{S})^{2}}{S-1}$. We simply repeated the estimation procedure of Oikarinen et al. [2023] that is summarized in Section 3.1 multiple times with different seed and data split to sample different activation vectors.
Distribution Fit. Inspired by ProbCBM proposed in Kim et al. [2023a], we estimate uncertainty from the data as a learnable parameter through distribution fitting. We assume a normal distrbution of the noise and model the standard deviation as a linear projection of the feature vector. The model is summarized below.
We obtain the observed score of a concept given an example: $\mathbf{x}$ using the multi-modal model: $g,g_{text}$ and the text description of the $k^{th}$ concept $T_{k}$. The concept score is modeled to be distributed by a normal distribution whose mean and standard deviation are linear functions of the feature representation of the model-to-be-explained: $f^{[-1]}(\mathbf{x})$. We optimize the value for $[\vec{p_{1}},\vec{p_{2}},\dots,\vec{p_{K}}],[\vec{q_{1}},\vec{q_{2}},\dots,\vec{q_{K}}]$ through gradient descent on the objective $\mathcal{\beta}=MLL(\mathcal{D},g)+\beta\times\mathbb{E}_{\mathcal{D}}\mathbb{E}_{k}[KL(\mathcal{N}(0,I)\|\mathcal{N}(\mu_{k}(\mathbf{x}),\sigma_{k}(\mathbf{x})))]$ very similar to the proposal of Kim et al. [2023a]. We picked the best value of $\beta$ and obtained $\epsilon^{DF}$ by averaging over all the examples: $\epsilon^{DF}=\mathbb{E}_{\mathcal{D}}[\sigma_{k}(\mathbf{x})]$.
Evaluation of different uncertainties.
We conduct our study using the ResNet-18 model pretrained on Places365 and Pascal dataset that were discussed in Section 6.1. We use human-provided concept annotations to train per-concept linear classifier on the representation layer. We retained only 215 concepts of the total 720 concepts that have at least two positive examples in the dataset. We then evaluated the per-concept linear classifier on a held-out test set to obtain macro-averaged accuracy. The concepts with poor accuracy are the ones that cannot be classified linearly using the representations. Therefore the error rate per concept is the ground-truth for uncertainty that we wish to quantify.
We may now evaluate the goodness of uncertainty: $\epsilon$ of U-ACE by comparing it with ground-truth (error-rate); observe they are both K-dimensional vectors. We report two measures of similarity in Table 9: (1) Cosine-Similarity (Cos-Sim) between $\epsilon$ and error-rate, (2) Jaccard Similarity (JS) (https://en.wikipedia.org/wiki/Jaccard_index) between top-k least uncertain concepts identified using error-rate and $\epsilon$. For any two vectors $u,v$, and their top-k sets: $S_{1}(u),S_{2}(v)$, the Cos-Sim and JS are evaluated as follows.
@@
Table 9: Evaluation of uncertainties estimated using U-ACE, MC sampling and Distribution Fit (see text for their description). Cos-Sim is the cosine-similarity with ground-truth value of uncertainty. The next three columns show Jaccard similarity between the top-k concepts ranked by ground-truth uncertainty and each of the three methods. Higher the better for all the values.
@@
H.2 Bayesian Estimation and Significance of Prior
@@
We replicate the study on Broden dataset of Section 6.1 of Table 1 with two new baselines. We replace the linear model estimation of U-ACE described in Section 3 with an out-of-the-box Bayesian Regression estimator available from sklearn [^1] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.BayesianRidge.html, which we refer as Bayes Regr. Effectively, Bayes Regr. is different from U-ACE only in the prior. We also compare with the estimation of fitting using Bayes Regr. but when the input is perturbed with the noise estimated by U-ACE. We refer to this baseline as Bayes Regr. with MC.
@@
H.3 Effect of Regularization Strength on Y-CBM and O-CBM
@@
Hyperparameter tuning is tricky for concept explanations since they lack a ground-truth or validation set. The reported results for Y-CBM and O-CBM in the main paper used the default value of the regularization strength of the corresponding estimator, which is Lasso [^2] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.Lasso.html for Y-CBM and LogisticRegression [^3] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.LogisticRegression.html for O-CBM. Both the estimators are part of sklearn. We had to reduce the default regularization strength of Y-CBM to $\alpha=10^{-3}$ so that estimated weights are not all 0. The $\kappa$ of U-ACE is somewhat arbitrarily set to $0.02$ on Broden dataset for sparse explanation with non-zero weight for only 20-30% of the concepts.
@@ Appendix I Evaluation using CUB dataset
Wah et al. [2011] released a bird dataset called CUB with 11,788 images and 200 bird species. Moreover, each bird image is annotated with one of 312 binary attributes indicating the presence or absence of a bird feature. Koh et al. [2020] popularized an improved version of the dataset that retained only 112 clean attribute annotations. We evaluate using the cleaner dataset released by Koh et al. [2020] owing to their popularity in evaluating CBMs. We train a pretrained ResNet-18 model using the training split of CUB dataset. We then compute the explanation (i.e. saliency of concepts) using the test split. Similar to the evaluation of Section 6.1, we quantify the quality of explanations using distance from explanations computed using true concept annotations when using Oracle.
@@
Table 12: Distance of top-k salient concepts computed using Oracle and different estimation methods shown in the first row (lower the better). TCAV does well overall and U-ACE performs the best among methods without access to concept annotations.
@@
Table 13: Average overlap between top-k salient concepts computed using Oracle and different estimation methods shown in the first row (higher the better). TCAV does well overall and U-ACE performs the best among methods without access to concept annotations.